OpenStudy (anonymous):

A fire truck is called to a scene. Three minutes later, a second truck is called. The first truck averages only 30 mi/h, but the second averages 60 mi/h. The trucks travel a total of 12 miles and arrive at the same time. How long from the first call did the trucks take to arrive? How far did each travel?

OpenStudy (anonymous):

first truck traveled 3 miles and second traveled 9 miles First truck took 6 mins and second truck 12 mins from the first call

OpenStudy (anonymous):

just solve the two equations x+y=12 60*(y/60-x/30)=3

OpenStudy (anonymous):

you got caroline??

OpenStudy (ybarrap):

\(V_1\) is the speed of the 1st truck \(V_2\) is the speed of the 1st truck \(t_1\) is the time of travel for the 1st truck \(t_2\) is the time of travel for the 2nd truck \(t_2=t_1+3\) The distance the 1st truck travels is \(V_1\times t_1\) The distance the 2nd truck travels is \(V_2\times t_2\) The total distance they travel is \(V_1\times t_1+V_2\times t_2=12\) $$ V_1 t_1+V_2 t_2=12\\ V_1t_1+V_2 (t_1+3)=12\\ V_1t_1+V_2 t_1=12-3V_2\\ \implies t_1(V_1+V_2)=12 -3V_2\\ t_1={{12 -3V_2} \over V_1+V_2} $$ You now have \(t_1\), from which you can get \(t_2\) and the distances traveled by each truck.